OpenStudy (anonymous):

@ganeshie8 Need help with Combinations

OpenStudy (anonymous):

If \[\left(\begin{matrix}n \\ 12\end{matrix}\right)=\left(\begin{matrix}n \\ 9\end{matrix}\right)\] find \[\left(\begin{matrix}n \\ 17\end{matrix}\right)\] and \[\left(\begin{matrix}22 \\ n\end{matrix}\right)\]

OpenStudy (anonymous):

\[*\left(\begin{matrix}n \\ 12\end{matrix}\right)=\left(\begin{matrix}n \\ 8\end{matrix}\right)\]

OpenStudy (anonymous):

Do you know about the symmetry of the binomial coefficient/combination formula?

OpenStudy (anonymous):

uh no

OpenStudy (anonymous):

maybe be i know but i dont know what its called

OpenStudy (anonymous):

It's easily seen in the Pascal triangle: \[\begin{array}{c|cccccc} n\backslash k&0&1&2&3&4&\cdots\\ \hline 0&1\\ 1&1&1\\ 2&1&2&1\\ 3&1&3&3&1\\ 4&1&4&6&4&1\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\ddots \end{array}\]

OpenStudy (anonymous):

If it's not clear, the the value in the table is \(\dbinom nk\), the row an the top indicates \(n\) and the column to the left is \(k\).

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

For example, when \(n=3\), you get \(\dbinom 31=\dbinom 32=3\). This is what I mean by symmetry.

OpenStudy (anonymous):

oh okay yea i know that pattern The numbers on the left side have identical matching numbers on the right side, like a mirror image.

OpenStudy (anonymous):

Right. So given that \(\dbinom n{12}=\dbinom n9\), you have enough info to determine the value of \(n\).

OpenStudy (anonymous):

my bad that should (n8) not (n9)

OpenStudy (anonymous):

\[\left(\begin{matrix}n \\ 12\end{matrix}\right)=\left(\begin{matrix}n \\ 8\end{matrix}\right)\]

OpenStudy (anonymous):

Right, I just saw that... Anyway, knowing that \(\dbinom n{12}=\dbinom n8\), we then know that \(\dbinom n{11}=\dbinom n{9}\), then the middle number of that row is given by \(\dbinom n{10}\) (the bottom numbers close in on a midpoint, so to speak).

OpenStudy (anonymous):

ok, yea

OpenStudy (anonymous):

From the table you should be able to see that when \(n\) is even, you get that one center value that's not repeated. For example, with \(n=2\), that center value is \(\dbinom 21=2\), and for \(n=4\) the center is \(\dbinom 42=6\). In general, then, if \(n=2k\) (i.e. \(n\) is even), we can say that \(\dbinom{2k}k\) is the center value. For this problem, you have the center \(\dbinom n{10}=\dbinom{2k}k\) which means \(k=10\), and so \(n=2k=20\).

OpenStudy (anonymous):

To check: http://www.wolframalpha.com/input/?i=20+choose+12+%3D+20+choose+8

OpenStudy (anonymous):

It was that simple?

OpenStudy (anonymous):

Yep, it's mostly pattern recognition. Don't forget to evaluate the other coefficients.

OpenStudy (anonymous):

wait how do you know 10 was a midpoint?

OpenStudy (anonymous):

That conclusion was based on the pattern for other even values of \(n\).

OpenStudy (anonymous):

okay, I understand on how to this now. Thanks :D

OpenStudy (anonymous):

yw